Project

General

Profile

2016-04-11.html

Tia Miceli, 10/31/2016 01:12 PM

 
1
<?xml version="1.0" encoding="UTF-8" standalone="no"?>
2
<!DOCTYPE html PUBLIC "-//W3C//DTD XHTML 1.0 Transitional//EN" "http://www.w3.org/TR/xhtml1/DTD/xhtml1-transitional.dtd">
3
<html><head><meta http-equiv="Content-Type" content="text/html; charset=UTF-8"/><meta name="exporter-version" content="Evernote Mac 6.10 (454267)"/><meta name="altitude" content="229.3861236572266"/><meta name="author" content="tiamiceli@gmail.com"/><meta name="created" content="2016-04-11 16:09:06 +0000"/><meta name="latitude" content="41.83864615109475"/><meta name="longitude" content="-88.26138208502648"/><meta name="source" content="desktop.mac"/><meta name="updated" content="2016-10-31 17:59:37 +0000"/><title>2016-04-11</title></head><body>
4
<div>Booster: (Bill)</div>
5
<ul>
6
<li>15 Hz ops. TLG changes to do 5 hours continuous at 15 Hz. Demonstrated for future.</li>
7
<li>recycler kickers have long rise time, need 3.</li>
8
<li>our trigger is set from first bunch filled, so not an issue. We would see one less bunch per spill if we were sensitive.</li>
9
<li>April 6, measurements on cable lengths.</li>
10
</ul>
11
<div><br/></div>
12
<div>BNB: (Tom)</div>
13
<ul>
14
<li>3 bunch notch from a 2 bunch notch</li>
15
<li>fast trim magnent is identified and tested</li>
16
<li>hardening on seal for horn 4 being tested if it’s just seal, or throughout bulk</li>
17
<li>Air fan replaced in target hall. (Backfish)</li>
18
<li>Dedicated collimator studies next week.</li>
19
</ul>
20
<div><br/></div>
21
<div>ANNIE: (Carrie)</div>
22
<ul>
23
<li>Last week commissioning HV system.</li>
24
<li>Fill tank w/ water this week Wednesday, ~8 hours.</li>
25
</ul>
26
<div><br/></div>
27
<div>MicroBooNE: (Andy)</div>
28
<ul>
29
<li>Downtime last week: test noise filter on service board Wed-Thur.
30
<ul>
31
<li>Factor x2 reduction in noise.</li>
32
<li>Study compression ratios of data &amp; implications for physics.</li>
33
<li>Issues w/ Sunday night SlowMon computer: RAID array firmware is updated. Issue went away last week.</li>
34
</ul>
35
</li>
36
<li>Full filter used instead of reduced one because reduced one was not filtering enough oxygen from the top-off argon.</li>
37
<li>Spare pump put into use, old pump was used for ~20% longer w/out maintenance.</li>
38
<li>Have weird # of triggers per spill</li>
39
</ul>
40
<div><br/></div>
41
<div><br/></div>
42
</body></html>